. Every minute, 5 x 10^19 electrons pass a particular cross section of wire.
a} Show the complete calculation of the amount of charge moving through the wire in
a minute
b) Show your complete calculation of the electric current in the wire.
1
Expert's answer
2020-02-10T09:43:51-0500

As per the given question,

time =1 min = 60 sec

a)

Total number electrons are passing through a cross section of wire = "5 \times 10^{19}" electrons

We know that , "Q=ne"

Where e is the charge on the electrons,

"e=1.6\times 10^{-19}C"

So, charge "Q=ne=5\times 10^{19}\times1.6\times10^{-19}=8C"

b)

We know that electric current "i=\dfrac{Q}{t}"

"\Rightarrow i=\dfrac{8C}{60sec}=\dfrac{2}{15}A=0.133A"